# A Novel Hybrid Time-Varying Graph Neural Network For Traffic Flow Forecasting

### v3

@@ V-G1 Sensitivity to Layers
The impact of the model’s layer count is illustrated in Figure 6(a). However, our findings indicate that increasing the number of layers does not yield a significant improvement in performance. This observation suggests that the HTVGNN model may have reached its predictive error limit.
@@ V-G2 Sensitivity to Heads
The impact of the number of heads on RTP multi-head attention is illustrated in Figure 6(b). It can be observed from the figure that while the attention score can be effectively computed with two heads, the dimensionality of the corresponding space for this head count is insufficient. As the number of heads increases, spatio-temporal information is mapped to a higher-dimensional space, resulting in more reasonable the attention score. However, an excessively large number of heads may introduce numerous negative high-dimensional spaces, which hinders accurate calculation of the attention score.
@@ V-G3 Sensitivity to Dimensions
The impact of the hidden embedding dimension D on the model HTVGNN is illustrated in Figure 6(c). As depicted, smaller dimensions fail to effectively encode spatiotemporal information, while excessively large ones may lead to overfitting.
@@ V-I Efficiency Study
Additionally, we conducted a comparison of the training and inference efficiency of HTVGNN, Bi-STAT, and PDFormer on the PEMS08 dataset. To ensure fairness in our evaluation, all methods were executed on a server equipped with an Intel Core i7 13700KF processor and a single NVIDIA RTX3090 graphics card. The results of this comparison are presented in Table X.
TABLE X: Combination Time on the PEMS04 and PEMS08
@@
As shown in Table X, despite the time-consuming nature of RNNs, the HTVGNN model utilizes a single layer. In contrast, PDFormer and Bi-STAT solely rely on transformer-based architectures that inherently support parallel processing. However, to capture deep spatio-temporal correlations, transformer-based models often require multiple layers; for instance, PDFormer employs six layers while Bi-STAT uses two layers. Nevertheless, due to the inclusion of a recollection process, this model incurs twice the original time complexity. Consequently, these transformer-only models do not offer any advantage in terms of time complexity. Conversely, our proposed HTVGNN model achieves superior prediction error performance on PEMS08 dataset with just one layer despite employing computationally intensive RNNs. Moreover, HTVGNN significantly outperforms PDFormer and Bi-STAT in prediction accuracy while maintaining similar computational requirements.
@@ VI Conclusion
In this paper, we propose a novel hybrid time-varying graph neural network prediction method, HTVGNN. To accurately capture the most relevant temporal features in temporal correlation modeling, we introduce a time-aware multi-head attention mechanism based on time-varying mask enhancement. We allocate the optimal attention calculation scheme for different temporal features. In spatial correlation modeling, we employ two strategies. Firstly, in static spatial correlation modeling, we design a coupled graph learning mechanism to integrate the graph learned at each time step and effectively enhance the model’s long-term prediction capability. Secondly, in dynamic spatial correlation modeling within the road network, we define a mask based on topological matrix and semantic matrix to optimize the dynamic time-varying graph and improve the model’s ability to capture dynamic spatial correlations. Simulation experiments conducted on four real datasets validate our proposed method HTVGNN by demonstrating its superior prediction accuracy compared with state-of-the-art spatio-temporal graph neural network models. Furthermore, our proposed coupling graph learning method significantly improves long-term prediction performance of the model as observed in ablation experiments. Future work will focus on exploring more effective masking techniques through model pre-training and providing more efficient yet powerful masking mechanisms for Transformer models.

### v4

@@ V-G1 Sensitivity to Layers
The impact of the model’s layer count is illustrated in Figure 6(a). Our findings indicate that increasing the number of layers does not yield a significant improvement in performance. This observation suggests that the HTVGNN model may have reached its predictive error limit.
@@ V-G2 Sensitivity to Heads
The impact of the number of heads on enhanced temporal perception multi-head self-attention is illustrated in Figure 6(b). It can be observed that the attention score calculated by the two heads can obtain a better prediction result. As the increase of the number of heads, spatio-temporal features will be mapped to a higher-dimensional space, resulting in more reasonable the attention score. However, an excessively large number of heads may introduce numerous negative interference and hinder accurate calculation of the attention score.
@@ V-G3 Sensitivity to Dimensions
The impact of the model embedding dimension $D$ on the model HTVGNN is illustrated in Figure 6(c). It can be seen that a smaller $D$ fails to effectively encode spatio-temporal features, while an excessively large $D$ may lead to overfitting.
@@ V-I Efficiency Study
Additionally, we conducted a comparison of the training and inference efficiency of HTVGNN, Bi-STAT, and STWave on the PEMS08 dataset. To ensure fairness in our evaluation, all methods were executed on a server equipped with an Intel Core i7 13700KF processor and a single NVIDIA RTX3090 graphics card. The results of this comparison are presented in Table X.
TABLE X: Combination Time on the PEMS08
@@
As shown in Table X, despite the time-consuming nature of RNNs, the HTVGNN model utilizes a single layer. In contrast, STWave and Bi-STAT solely rely on transformer-based architectures that inherently support parallel processing. However, to capture deep spatio-temporal correlations, transformer-based models often require multiple layers. For instance, the layers of both Bi-STAT and STWave are two. Nevertheless, due to the facts that Bi-STAT includes a recollection process and STWave includes a dual-channel spatio-temporal network to model trends and events, the complexity of these two models is high. Consequently, these transformer-only models do not offer any advantage in terms of time complexity. In comparison, although the computationally intensive RNNs was employed into our proposed one layer HTVGNN model, HTVGNN achieves a superior prediction performance on PEMS08 dataset. Moreover, HTVGNN significantly outperforms STWave and Bi-STAT in prediction accuracy while maintaining similar computational costs.
@@ VI Conclusion
In this paper, we propose a novel hybrid time-varying graph neural network prediction method, HTVGNN. To accurately capture the most relevant temporal features in temporal correlation modeling, we introduce an enhanced temporal perception multi-head self-attention mechanism based on time-varying mask enhancement. We allocate the optimal attention calculation scheme for different temporal features. In spatial correlation modeling, we employ two strategies. Firstly, in static spatial correlation modeling, we design a coupled graph learning mechanism to integrate the graph learned at each time step and effectively enhance the model’s long-term prediction capability. Secondly, in dynamic spatial correlation modeling within the road network, we define a mask based on topological matrix and semantic matrix to optimize the dynamic time-varying graph and improve the model’s ability to capture dynamic spatial correlations. Simulation experiments conducted on four real datasets validate our proposed method HTVGNN by demonstrating its superior prediction accuracy compared with state-of-the-art spatio-temporal graph neural network models. Furthermore, our proposed coupling graph learning method significantly improves long-term prediction performance of the model as observed in ablation experiments. Future work will focus on exploring more effective masking techniques through model pre-training and providing more efficient yet powerful masking mechanisms for Transformer models.